### v2

@@ 3 Challenges and Opportunities
3.0.1 Ethical Issues
Biases present in training datasets can lead to skewed results, disproportionately affecting certain patient groups [195]. Future efforts must be directed towards developing ethical frameworks and strategies that mitigate biases, guaranteeing equitable applications of LLMs in healthcare.
It is essential to ensure patient autonomy, data confidentiality, and protection against breaches. The development of domain-specific, open-source models for on-premises use is a notable advancement in preserving patient privacy and reducing dependency on external entities. Further specialization of models and adoption of computationally efficient solutions can ensure controlled data access and compliance with local policies on data handling.
3.0.2 Datasets
@@
3.0.3 Factuality Challenges
The issue of factuality is critical in healthcare. Medical solutions need to have access to up-to-date and proven sources and must also undergo rigorous checks for factual accuracy. There is an urgent need to develop automated metrics for assessing the factual accuracy of medical text generation [196]. Moreover, research exploring the integration of medical LLMs with ontologies, graph attention networks, and other more deterministic models represents another promising direction.
3.0.4 Human-AI Collaboration
@@
3.0.5 Necessity for Empirical Studies
@@
4 Summary and Conclusions
This study offers a detailed exploration of recent progress in large language models in the medical field, with a special focus on clinical applications. It begins by tracing the evolution of LLMs, covering both general-purpose and domain-specific models, and examines their architectures and medical applications. The paper highlights various tasks these models perform, such as text generation, token classification, and question answering, illustrating their utility with real healthcare scenarios. This demonstrates how LLMs enhance both efficiency and effectiveness in medical practice, particularly through the use of open-source models that support custom solutions and protect privacy.
@@ 4 Summary and Conclusions
Recent advancements in LLM research have introduced more comprehensive and efficient approaches, particularly through multimodal models that amalgamate visual, textual, and potentially audio data. These advancements facilitate holistic AI solutions and are complemented by techniques like parameter-efficient fine-tuning and flash attention, which minimize computational demands. The shift towards generative LLMs with in-context learning capabilities marks a critical evolution, specifically within niche medical areas such as radiology report generation.
However, the deployment of LLMs in healthcare presents several challenges. Key among these are maintaining patient autonomy, ensuring data confidentiality, and safeguarding against data breaches. The application of open-source, domain-specific models directly within healthcare facilities can mitigate these issues while reducing dependency on third-party data processors. Additionally, training data biases that could affect patient groups unequally need addressing through comprehensive ethical frameworks designed for equitable healthcare.
To improve factual accuracy, ongoing access to updated sources and robust validation metrics are essential. Integrating LLMs with established medical ontologies, deterministic models, or structured data can enhance their utility. Moreover, fostering human-AI collaborations can improve clinical decision-making and alleviate routine burdens, potentially reducing professional burnout.
Future research should focus on empirical studies in real-world settings to better understand how AI can be integrated into healthcare workflows, its impact on patient care, and its long-term effects on outcomes and costs. This will address practical challenges and refine the deployment strategies for LLMs in healthcare, ensuring they meet the dual goals of efficacy and ethical integrity.

### v3

@@ 4 Discussion
4.1 Evaluation Challenges
The evaluation of language models in the medical domain is a multifaceted challenge. One major issue arises from the technical complexity of assessing model performance on tasks with minimal human supervision. For instance, while classification tasks benefit from well-established metrics such as accuracy, precision, recall, and F-measure, evaluating models on more complex tasks, such as medical conversations, remains technically challenging. These tasks often require human assessment or intricate evaluation frameworks [185, 186, 225, 226].
Another challenge arises from unique domain-specific issues, where standard model evaluation from a purely technical perspective is often insufficient. Abbasian et al. [227] categorize evaluation metrics into intrinsic metrics, such as Perplexity, which measure internal language proficiency and coherence, and extrinsic metrics, which assess real-world impact and the model’s ability to meet human-centric expectations. Within extrinsic metrics, the authors identify general metrics applicable to all tasks, such as trustfulness, bias, and toxicity, as well as domain-specific metrics, such as up-to-dateness and empathy. While the intrinsic/extrinsic categorization provides a useful theoretical framework for understanding evaluation metrics, in practice, metrics are often organized based on their functional role in model evaluation. Table 5 presents a task-oriented categorization, where metrics are grouped by their evaluation purpose.
{adjustwidth}
-2.25in0in
Table 5: Evaluation Metrics for Major Tasks in Medical Domain using Language Models
Tasks Task-Specific Metrics Extrinsic Generic Metrics and Benchmarks Extrinsic Domain-Specific Metrics and Benchmarks
Classification and Information Extraction: Clinical Acronym Disambiguation, Eponyms Disambiguation, Phenotyping, Medical Coding, Modeling Patient Timeline, Social Media Monitoring, Querying Data from Electronic Health Records, Extracting Information from Clinical Narrative Reports, Extracting Information From Medical Articles Accuracy, Precision, Recall, F-Score, AUC Necessary Only When Using Generative Models: Metrics for evaluating LLM performance in general (e.g., Academic Benchmarks, Factuality, Complex Reasoning) [228, 5]. Privacy [229] (when patient data is involved)
Generation and Summarization: Medical Report Generation, Clinical Note Generation, Generating Summaries For Laypersons, Generating Summaries for Patient-Provider Dialogues, Generating Textual Descriptions From Graph Models, Summarizing Clinical Study Reports, Summarizing Patient-Provider Dialogues, Simplification of Medical Texts, Simplification of Radiology Reports, Improving Biomedical Text Readability BLEU, ROUGE, METEOR, Perplexity [14], BERTScore Metrics for evaluating LLM performance in general (e.g., Academic Benchmarks, Factuality, Complex Reasoning). Reliability [227], Up-to-dateness [230], Privacy
System-Patient Conversation: Mental Health Bots, Medical Chatbots and Health Assistants Match Rate, Dialogue Accuracy, Average Request Turn, Complex Chatbot-Specific Metrics (e.g., DEAM [226], DynaEval [225]) Metrics for evaluating LLM performance in general (e.g., Academic Benchmarks, Factuality, Complex Reasoning), Safety and Bias [231, 232, 227, 233, 234]. Reliability, Up-to-dateness, Privacy, Empathy [227]
Unlike task-specific metrics, which are widely adopted and often have clear mathematical formulations, extrinsic metrics involving evaluation of LLMs are significantly more technically complex due to their reliance on specialized frameworks, human alignment, or external models. For example, DynaEval [225] is a framework-based metric that employs a graph convolutional network to evaluate model performance. Benchmarks designed for specific goals, such as TruthfulQA [234], RealToxicityPrompts [233], and BBQ [232], may fail to generalize well to multilingual tasks or domain-specific applications due to their reliance on carefully crafted datasets tailored to specific scenarios. Furthermore, some extrinsic metrics rely on external models for evaluation [226], introducing stochasticity into the process due to the non-deterministic nature of model outputs.
Moreover, the probabilistic nature of LLMs makes it challenging to evaluate their performance with respect to domain-specific metrics, especially for unseen input combinations or highly specialized, out-of-distribution data.
Notably, the demand for domain-specific extrinsic metrics grows as models are opened to a broader audience performing more complex tasks. In the healthcare sector, domain-specific evaluation is particularly critical due to its unique ethical and regulatory requirements. Hond et al. [235] highlight the importance of complementing general and domain-specific evaluation with clinical impact validation, a process that assesses outcomes such as improved health results, higher patient satisfaction, or reduced administrative burdens. However, to the best of our knowledge, no existing frameworks can automatically perform clinical impact validation with reasonable accuracy.
4.2 Ethical Issues
The application of language models in clinical settings raises substantial ethical concerns. To address these, foundational principles such as respect of persons, beneficence, and justice, as outlined in the Belmont Report [236], provide a guiding framework. Solomonides et al. [237] further expand this by emphasizing technical principles like fairness, interpretability, and explainability, alongside organizational principles such as transparency, accountability, and benevolence. While these principles offer a comprehensive ethical blueprint, implementing them effectively in real-world systems remains a considerable challenge. To navigate these challenges, we propose a tiered approach that establishes progressively stricter levels of compliance with ethical principles. Each level represents a set of actionable system properties derived from ethical principles, ensuring they are translated into concrete requirements. Systems must satisfy the foundational properties of lower levels before progressing to higher ones. This structure enables a practical approach to balancing ethical integrity with technological feasibility while accommodating gradual improvements in system capabilities.
Level 1 establishes a baseline set of system properties required for safe, minimal-risk usage. Given the inherent risk of generating misleading or factually incorrect content [177, 187, 217], the first Level 1 requirement focuses on the principle of nonmaleficence. Additionally, ensuring the confidentiality and security of patient data is critical, as emphasized by Ong et al. [238], to uphold patient autonomy and system dependability, both central to Level 1 requirements. These safeguards can be achieved through strict data management protocols and local deployments that minimize data leakage. Robust anonymization techniques [179] play a key role in securing patient privacy when using data externally.
Systems that comply with this level may be acceptable in lower-risk contexts, where errors can be quickly identified and corrected by qualified professionals. Adhering to Level 1 properties enables the ethical use of language models for a variety of tasks, including most classification and information extraction tasks, as well as many summarization and paraphrasing tasks.
Level 2 builds upon Level 1 compliance and introduces additional principles: fairness, interpretability, auditability, and knowledge management. These enhanced requirements enable systems to perform more sensitive tasks, such as decision support, while avoiding direct system-patient interactions. However, modern LLMs often lack interpretability and remain susceptible to bias [239, 240, 241], a challenge further exacerbated by the absence of well-established, widely applicable metrics to measure these properties. This makes the integration of Level 2 principles an open research area. As a result, the deployment of LLMs in high-stakes applications, such as differential diagnosis, triaging, or modeling patient timelines, should generally involve human-in-the-loop supervision to mitigate risks.
Level 3 encompasses the principles of beneficence, justice, explainability, and benevolence, which are critical for systems capable of interacting with patients without supervision from healthcare professionals. A system can achieve Level 3 only after fully satisfying the requirements of Levels 1 and 2. Attaining this tier would enable widespread, ethically responsible deployment across all reviewed medical tasks. However, to the best of our knowledge, no current models or systems meet this standard. Table 6 summarizes the proposed hierarchical levels, outlines scenarios of use for models and systems that comply with the corresponding ethical principles adopted from Solomonides et al. [237], and provides representative tasks for each level.
{adjustwidth}
-2.25in0in
Table 6: Levels of Ethical Compliance and Corresponding Scenarios in System or Model Applications
Level Ethical Principles Compliance Scenarios of Use Sample Tasks
Level 1 Nonmaleficence, Autonomy, Dependability - Non-critical tasks with low risk to patients
- Tasks under supervision of professionals
- Tasks that exclude direct interaction of patients with a system
Phenotyping, Medical Coding, Eponyms Disambiguation, Clinical Acronym Disambiguation, Generating Textual Descriptions from Graph Models, Generating Summaries for Patient-Provider Dialogues, Clinical Note Generation, Medical Report Generation, Querying Data from Electronic Health Records, Extracting Information from Clinical Narrative Reports, Extracting Information from Medical Articles, Summarizing Clinical Study Reports, Summarizing Patient-Provider Dialogues, Anonymization of Medical Documents
Level 2 Fairness, Interpretability, Auditability, Knowledge Management, Accountability - Tasks that involve decision-making without direct interaction of patients with a system
Triaging, Differential Diagnosis, Social Media Monitoring, Modeling Patient Timeline
Level 3 Beneficience, Justice, Explainability, Benevolence - Patient-facing applications Medical Chatbots and Health Assistants
From an ethical standpoint, current language model systems are far from fully ready for deployment in high-stakes clinical settings. While many systems can meet the baseline requirements of Level 1 for low-risk tasks under professional supervision, substantial challenges remain in achieving compliance with higher levels. The lack of widely adopted interpretability, fairness, and accountability metrics hampers progress toward Level 2. To our best knowledge, no existing systems currently meet the stringent requirements of Level 3, which are essential for patient-facing applications.
4.3 Datasets
@@
4.4 Human-AI Collaboration
@@
4.5 Necessity for Empirical Studies
@@
5 Conclusion
This study provides an in-depth examination of recent advancements in language models within the medical domain, with a particular emphasis on clinical applications and locally deployable solutions. It traces the development of language models, exploring both general-purpose and domain-specific architectures, and evaluates their role in medical contexts. The study highlights key tasks performed by these models, such as text generation, token classification, and question answering, demonstrating their practical utility through real-world healthcare scenarios.
@@ 5 Conclusion
Recent advancements in the field have introduced more comprehensive approaches, particularly through multimodal models that seamlessly integrate visual and textual data. These innovations enable holistic AI solutions and are bolstered by techniques such as parameter-efficient fine-tuning and flash attention, which significantly reduce computational requirements. The rise of generative LLMs with in-context learning capabilities marks a pivotal evolution, unlocking new possibilities in specialized medical domains like radiology report generation and medical chatbots - tasks that were considered unattainable just a decade ago.
However, deploying language models in healthcare presents several challenges. The first challenge lies in evaluating generative models, especially for complex tasks like medical conversations or summarizations, where task-specific metrics are insufficient, and extrinsic evaluations require intricate frameworks or human alignment. Additionally, critical domain-specific requirements, such as privacy, up-to-dateness, and empathy, are difficult to quantify or standardize in healthcare applications. Lastly, the lack of automated frameworks for clinical impact validation complicates the evaluation process further, hindering the ability to assess real-world outcomes, such as improved patient care or administrative efficiency.
The second challenge revolves around the ethical deployment of language models in clinical settings. While many systems meet baseline compliance for low-risk tasks under professional supervision, advancing to higher levels of ethical compliance remains demanding. Requirements such as fairness and interpretability are hindered by the absence of widely adopted metrics and the persistent biases in modern LLMs. Fully satisfying the comprehensive set of ethical standards essential for patient-facing applications is particularly daunting, and, to the best of our knowledge, no current systems meet these stringent requirements.
Mitigating these challenges requires both technological advancements and focused research. Data privacy concerns can be addressed through locally deployed models or robust anonymization techniques, while the up-to-dateness requirement can be managed using retrieval-augmented generation techniques for generative models or continuous adaptation for encoder-based models. These advancements make a wide range of tasks, those that exclude direct patient interaction and unsupervised decision-making, technically viable for broader adoption within medical organizations. However, even these tasks may require additional regulatory approvals and evaluation in real-world clinical settings before widespread implementation [242].
For tasks involving unsupervised decision-making, addressing privacy and model up-to-dateness alone may not suffice. This is especially true for generative models producing unstructured text, as there is a lack of widely adopted automated metrics for assessing both ethical compliance and performance. A practical approach for handling such tasks is to reformulate them into well-defined tasks, such as classification or information extraction, along with incorporating the latest developments in reasoning and interpretability [243, 244, 245]. In these contexts, established task-specific metrics are sufficient to evaluate performance, and mathematical formulations for assessing ethical components like fairness [209] are readily available. If the approach with reformulating the tasks is not feasible, research exploring the integration of LLMs with ontologies, graph attention networks, and other more deterministic and interpretable models represents another promising direction in moving the adoption of models feasible for unsupervised decision-making.
Finally, we find no existing solutions that are ethically and technically prepared for tasks involving direct interaction with patients beyond purely administrative functions.
In general, the specific demands of the medical field, where precision is paramount due to the high cost of errors and ethical compliance must address a wide range of principles such as interpretability, fairness, and accountability, result in an unavoidable temporal gap between technological advancements and their adoption. To address this gap, future research, besides purely technological advancements, should prioritize empirical studies in real-world settings to explore how AI can be seamlessly integrated into healthcare workflows without increasing the burden on healthcare professionals. Additionally, it should assess the impact of AI on patient care and its long-term implications for outcomes and costs. We also advocate for the development of Medical Model Cards, inspired by the generic Model Cards proposed by Mitchell et al. [246]. These Medical Model Cards should accompany all models intended for use in clinical settings, providing detailed information about compliance with ethical principles, validated through appropriate metrics or benchmarks. They should also include the intended tasks and corresponding performance benchmarks. This framework will facilitate quicker and more informed model selection for adoption in the healthcare domain.